### v2

# Toward Metaphor-Fluid Conversation Design for Voice User Interfaces

Abstract.
Metaphors play a critical role in shaping user experiences with Voice User Interfaces (VUIs), yet existing designs rely on static, human-centric metaphors that fail to adapt to diverse contexts and user needs. We introduce Metaphor-Fluid Design, a novel approach that adjusts metaphorical representations based on conversational use-contexts. We compare this to a Default VUI, which models commercial VUIs commonly designed around the assistant persona, offering uniform interaction across contexts. In Study 1 (N=130), metaphors were mapped to four key use-contexts—commands, information seeking, sociality, and error recovery—along dimensions of formality and hierarchy, revealing distinct preferences for task-specific metaphorical designs. Study 2 (N=91) evaluates a Metaphor-Fluid VUI against a Default VUI, showing that Metaphor-Fluid Design enhances perceived intention to adopt, enjoyment, and likability. However, individual differences in metaphor preferences highlight the need for personalization. These findings challenge the one-size-fits-all paradigm and demonstrate Metaphor-Fluid Design’s potential for more adaptive human-AI interactions.
Conversational Agents, Voice User Interfaces, Metaphors, Personas, Design
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Auditory feedbackccs: Human-centered computing Natural language interfacesccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
Figure 1. The schematic depicts the proposed evolution toward Metaphor-Fluid Design, starting from the description of Voice User Interfaces (VUIs) in different social roles using popular metaphors [A], leading to the identification of design issues [B], and ultimately inspiring Metaphor-Fluid design [C].
1. Introduction
Voice User Interfaces (VUIs) are becoming an integral part of daily interactions, embedded in smart speakers, mobile devices, automobiles, and wearable technologies. As these technologies proliferate, the ways in which users engage with and conceptualize them evolve dynamically. However, contemporary VUI design remains largely constrained by a single dominant metaphor—humanness (Doyle et al., 2019). Most commercial VUIs are designed to mimic human conversational abilities, using speech, turn-taking, and natural language processing to create an interaction model that feels intuitive (Edlund, 2019; Pradhan and Lazar, 2021). Within this overarching metaphor, VUIs are often framed as social entities that take on various roles, including teachers, therapists, coaches, advisors, and companions (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019; Trajkova and Martin-Hammond, 2020; Purington et al., 2017). Among these, the assistant persona has emerged as the most prevalent, dominating commercial implementations and positioning the VUI as a helpful, subservient entity designed to execute user commands (Desai and Twidale, 2023; Doyle et al., 2021). While this metaphor provides a familiar interaction paradigm, it also imposes significant limitations. It assumes a fixed, one-size-fits-all identity for the VUI, failing to account for the fluid or context-dependent nature of human-VUI interactions across different tasks and contexts (Pradhan and Lazar, 2021; Desai et al., 2024b; Desai and Chin, 2023; Desai and Twidale, 2022). This rigidity leads to usability challenges (Cowan et al., 2017), mismatches between user expectations and system behavior (Doyle et al., 2021), and broader concerns related to identity surrounding anthropomorphic design choices (McMillan and Jaber, 2021; Desai and Twidale, 2022; Simpson and Crone, 2022; Pradhan and Lazar, 2021; Luger and Sellen, 2016; Turk, 2016).
These challenges reflect a holistic issue in interface design: the role of metaphors in shaping user interaction. Metaphors have played a foundational role in shaping Human-Computer Interaction (HCI) (Neale and Carroll, 1997; Laurel, 1997; Laurel and Mountford, 1990; Carroll and Thomas, 1982; Carroll et al., 1988; Don et al., 1992). In Graphical User Interfaces (GUIs), they are structured into primary and secondary layers: the desktop metaphor, for instance, serves as the primary organizing concept, with secondary metaphors—such as files, folders, and recycling bins—providing finer interaction affordances (Blackwell, 2006). These layered metaphors help users form coherent mental models of system functionality. In contrast, VUIs lack a stable visual representation, making their metaphorical framing less explicit. Instead, this framing is operationalized through persona design, where a VUI’s voice, linguistic style, and response behaviors collectively establish its identity (Pradhan and Lazar, 2021). Unlike GUIs, where metaphors map interactions onto familiar objects, VUIs inherently imply humanness, reinforcing anthropomorphic interpretations. The presence of human-like speech cues, names, and conversational norms encourages users to assign social roles (e.g., therapist or companion) to VUIs, often expecting them to behave with intelligence, adaptability, and even emotional awareness (Luger and Sellen, 2016; Sutton et al., 2019).
Despite this anthropomorphic framing, users do not engage with VUIs as static entities (Pradhan et al., 2019). Instead, they shift their metaphorical framings depending on use-contexts—the functional settings in which interactions take place (Desai and Twidale, 2023). Prior research has identified several recurring use-contexts in VUI interactions, including commands, where users expect efficiency and compliance (e.g., ”Turn off the lights”); information seeking, where they seek structured knowledge retrieval (e.g., ”What is the capital of Japan?”); sociality, where they engage in casual conversation (e.g., ”Tell me a joke”); and error recovery, where they attempt to correct or refine system responses (e.g., ”That’s not what I meant—try again”) (Sciuto et al., 2018; Desai and Chin, 2023; Ammari et al., 2019; Bentley et al., 2018). In each of these use-contexts, users adopt distinct metaphorical descriptions: they conceptualize a VUI as a butler or secretary when issuing commands, as an expert librarian when retrieving information, and as a companion in social exchanges (Desai and Twidale, 2023). However, commercial VUIs impose a singular, unchanging persona across all contexts, failing to adapt to the shifting expectations users bring to different types of interactions. This rigidity results in fundamental misalignments—when users anticipate authoritative expertise but receive generic, non-committal responses, or when they expect a social exchange but encounter the same neutral, task-oriented demeanor, the interaction feels unnatural and disengaging (Desai et al., 2023a). These breakdowns not only present usability challenges but also raise deeper ethical concerns.
The assistant metaphor, in particular, has been widely critiqued for reinforcing problematic socio-cultural assumptions (McMillan and Jaber, 2021; Brahnam and De Angeli, 2008; Brahnam et al., 2011). Acting as the Default VUI in commercial systems such as Amazon Alexa, Google Assistant, and Apple Siri, this metaphor frames VUIs as subservient entities designed to efficiently execute user commands (Turk, 2016). These systems are marketed as digital assistants, emphasizing compliance, convenience, and deference—qualities that are often reinforced through feminized voices and polite, deferential speech patterns (Kuzminykh et al., 2020). This design choice perpetuates outdated power dynamics that position technology as a passive, compliant agent rather than an adaptive, context-aware tool. Prior studies have highlighted the risks of over-anthropomorphizing VUIs, where users may develop unrealistic expectations of intelligence or emotional understanding, leading to frustration when the system fails to respond appropriately (Cowan et al., 2017; Trajkova and Martin-Hammond, 2020). At the same time, users may dehumanize VUIs, treating them with excessive aggression or dismissiveness—behaviors that some researchers argue could spill over into human social interactions (Laurel, 1997; Brahnam and De Angeli, 2008). Furthermore, the universal application of a single persona disregards the diverse ways in which users naturally interpret and interact with these systems.
Building on prior work, we introduce and propose a shift toward Metaphor-Fluid Design (as illustrated in Figure 1), a novel design approach that moves beyond static persona models by enabling VUIs to dynamically shift their metaphorical framing based on conversational context. Rather than adhering to a singular metaphor—such as the assistant—this approach allows the system to adopt multiple personas, aligning its interaction style with the expectations of the task at hand. Crucially, Metaphor-Fluid Design also incorporates non-human and fictional metaphors, which have been shown to be both relevant and frequently used by users (Desai et al., 2024b; Desai and Twidale, 2023; Jung et al., 2022). While commercial VUIs default to anthropomorphic personas, users often describe them in non-human terms, likening them to machines, calculators, encyclopedias, or even mythical entities like genies or ghosts (Desai and Twidale, 2023). By broadening the metaphorical landscape beyond human roles, this approach creates greater flexibility in interaction design, allowing VUIs to respond more intuitively to diverse user expectations. This work is guided by two primary research questions:
  • RQ1: What metaphors commonly align with users’ interactions in use-contexts such as commands, information seeking, sociality, and error recovery during conversations with VUIs?
  • RQ2: How does a Metaphor-Fluid VUI compare to a Default VUI in terms of perceived enjoyment, intention to adopt, trust, likability, and intelligence, and in the characteristics of the metaphorical descriptions it generates?
These questions are explored through two studies. The first study (N=130) maps user-generated metaphors to different conversational contexts, revealing systematic variations in metaphorical preferences across tasks. Findings show that users do not conceptualize VUIs through a singular metaphor but instead shift their framings based on the use-context. The second study (N=91) empirically evaluates a Metaphor-Fluid VUI—which adapts its metaphorical framing dynamically—against a Default VUI that consistently operationalizes the assistant metaphor. Results indicate that Metaphor-Fluid design leads to significantly higher perceived likability, enjoyment, and intention to adopt, showing that aligning a VUI’s persona with context-specific expectations positively impacts user perceptions. However, findings also highlight individual differences in metaphor preferences.
Findings from this work challenge the dominant paradigm of static persona design in VUIs. By formalizing Metaphor-Fluid Design as a structured approach to VUI development, this paper contributes to a growing discourse on how metaphor selection influences user interaction. These insights have broader implications for the design of conversational agents, underscoring the importance of flexibility, contextual awareness, and ethical considerations in shaping future human-AI conversational interactions.
2. Related Work
VUIs enable users to interact naturally through spoken language and are embedded in a range of consumer technologies, including smart speakers, smartphones, headphones, and even household appliances. Popular examples include Amazon Alexa, Apple Siri, and Google Assistant, which have become integral to everyday interactions. However, the term VUI is often used interchangeably with other labels, such as Intelligent Personal Assistants, Digital Assistants, Conversational Agents (CAs), Conversational User Interfaces (CUIs), and Speech Interfaces (Clark et al., 2019). While these terms share similarities, they differ in scope. For instance, CAs and CUIs encompass both voice- and text-based systems, including chatbots, whereas VUIs specifically refer to interfaces that rely on voice as the primary input and output modality.
Given this terminological overlap, it is essential to clarify the scope of this research. This article focuses explicitly on disembodied, software-based VUIs—systems that process and generate spoken language without physical embodiments, such as robots or avatars. However, given the broader relevance of CUIs in conversational AI research, we reference CUIs where applicable to contextualize and discuss our findings within the wider discourse on human-AI interaction.
2.1. Metaphors—A Brief Primer
Metaphors are more than just literary flourishes; they are fundamental tools for understanding and interpreting the world (Indurkhya, 2013). In linguistics and cognitive science, metaphors have been shown to shape not only how we communicate but also how we think (Gentner and Hoyos, 2017). According to Lakoff and Johnson’s theory of conceptual metaphors, our understanding of abstract concepts is often structured by metaphors derived from concrete experiences (Lakoff and Johnson, 1980). For example, when we say ”time is money,” we are using a metaphor to conceptualize time as a valuable resource, which affects how we perceive and manage it (Lakoff and Johnson, 1980).
Metaphors are powerful because they influence cognition and behavior, often in subtle ways. They help bridge the gap between the familiar and the unfamiliar, making complex ideas more accessible (Gentner and Hoyos, 2017). By providing a conceptual structure, metaphors allow individuals to grasp abstract concepts by relating them to more concrete experiences. In cognitive psychology, metaphors have been found to shape reasoning, decision-making, and problem-solving processes, thereby playing an integral role in how knowledge is constructed and applied (Moser, 2000; Cameron and Maslen, 2010). Through metaphors, we can connect domains that are seemingly unrelated, fostering creativity and enabling us to generate novel ideas (Lockton et al., 2019). In science and technology, metaphors often serve as foundational frameworks that influence both how researchers theorize and how laypeople understand complex phenomena (Hofstadter, 1995). Alan Turing, for example, used a metaphor of a human computer following a set of instructions to describe his model for computation, which became the basis for modern computer science (Piccinini, 2003). Similarly, metaphors like ”genetic code” or ”cellular machinery” have shaped how scientists and the public conceptualize biological processes, providing familiar reference points that help make sense of otherwise abstract concepts (Keller, 2003). Metaphors thus provide a cognitive shortcut that allows individuals to comprehend and manipulate ideas beyond the immediate grasp of direct experience (Black, 1962).
Moreover, metaphors are deeply embedded in everyday language, often so much so that we are unaware of their influence. Phrases such as ”grasping an idea” or ”feeling down” illustrate how metaphorical thinking is woven into our conceptual system (Lakoff and Johnson, 1980).
2.2. Metaphors in HCI
Metaphors in HCI have profoundly shaped how users conceptualize and interact with technology. By leveraging familiar concepts (e.g., files, folders), metaphors bridge the gap between users’ prior knowledge and new digital environments, simplifying interactions and reducing cognitive load (Neale and Carroll, 1997; Carroll and Thomas, 1982). The desktop metaphor stands out as a pivotal example in GUIs, drawing on the physical analogy of an office desk to provide users with a coherent mental model for navigating digital spaces (Johnson-Laird, 1983). This foundational desktop metaphor is supplemented with secondary metaphors like folders, files, bins, scrolling, and icons and translates abstract computational processes into familiar actions, such as ”dragging a file to the trash” to discard a document (Colburn and Shute, 2008).
The impact of the desktop metaphor extends far beyond mere visual representation. It has played a crucial role in making computers more approachable and familiar, significantly contributing to their widespread adoption (Carroll et al., 1988). By aligning user expectations with system capabilities, the desktop metaphor has fostered increased usability and user satisfaction. Furthermore, it has provided a common language for developers and users to communicate about new technologies, facilitating the ongoing evolution of human-computer interfaces. The desktop metaphor’s influence on early computing systems was so profound that it shaped users’ mental models for decades. For a detailed exploration of its history and significance, Alan Blackwell’s work offers an in-depth analysis that underscores the metaphor’s lasting impact on HCI design principles (Blackwell, 2006).
2.3. Metaphors in VUIs
The study of metaphors in VUIs is relatively new compared to their rich history in HCI. The foundational ”humanness” metaphor serves as the principal metaphor in VUI design, where VUIs are often deployed in human-like social roles for VUIs, ranging from educators and coaches to storytellers and even therapists (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019). This approach aims to create a sense of familiarity by aligning VUIs with known social roles, thereby making interactions more intuitive and accessible (McMillan and Jaber, 2021; Desai and Twidale, 2023). VUIs like Amazon Alexa and Apple Siri are designed to function as digital assistants, drawing on the metaphor of the helpful human assistant to facilitate tasks such as playing music, providing information, managing smart home devices, and offering social companionship (Sciuto et al., 2018). These devices leverage the designer-intended and manufacturer-marketed (Turk, 2016) humanness metaphor to create an experience that feels natural and social, helping to bridge the gap between technology and everyday life (Olmstead, 2017; Auxier, 2019).
In practice, metaphors are operationalized in VUI design through system personas and form a crucial part of Conversation Design workflows (Sadek et al., 2023). Google’s Conversational UX guidelines [^1] https://developers.google.com/assistant/conversation-design/create-a-persona, for instance, emphasize role production by designing agents that mimic real-world figures, such as bank tellers or exercise coaches, situating these interfaces in familiar contexts (McMillan and Jaber, 2021). Designers often default to human roles because of the innate tendency to anthropomorphize technology. Anthropomorphism—the attribution of human traits to non-human entities—is central to VUI design, as it offers users a relatable interaction model (Epley et al., 2007).
The choice of metaphor in VUI design significantly impacts user perception. Studies have shown that metaphors signaling high competence initially attract users, while those suggesting lower competence tend to be favored in the long term (Khadpe et al., 2020). Furthermore, the formality of the VUI’s conversational style influences users’ metaphorical interpretations, with more formal styles evoking professional metaphors and informal styles eliciting personal ones (Chin et al., 2024). Research has also indicated a correlation between metaphor-projected personalities and user personalities, with users preferring metaphors that align with their own traits (Braun et al., 2019). Despite designers’ efforts to present clear mental models through personas, users often develop their own conceptual understandings of VUIs through folk theories (DeVito et al., 2018). These theories, often expressed as metaphors, attempt to explain VUI functionality or rationalize system errors (Kim and Choudhury, 2021; Kuzminykh et al., 2020; Desai and Twidale, 2022). The disparity between user expectations based on provided metaphors and actual VUI performance can lead to usage issues, as described by Norman’s ”gulf of execution” concept (Norman, 2013; Luger and Sellen, 2016).
Broadly, there are three significant challenges associated with the current implementation of metaphors in VUI design:
  1. (1)
    Ethical Issues: Designing VUIs based on roles like ”assistant” or ”servant” can inadvertently reinforce outdated stereotypes related to gender and power dynamics (McMillan and Jaber, 2021; Pradhan and Lazar, 2021). These metaphors may implicitly communicate roles of subservience, which can perpetuate harmful social norms, especially when the personas are given specific voices that align with stereotypical attributes. For example, the choice of a female voice for assistant roles often reinforces traditional gender stereotypes about caregiving and subservience, which has raised ethical concerns regarding the impact of such designs on societal attitudes (Brahnam et al., 2011; Turk, 2016; Kuzminykh et al., 2020).
  2. (2)
    Simplistic Representations: While users develop complex and context-dependent mental models of VUIs that change with the type of interaction or domain of use, designers often rely on a single overarching metaphor to represent these systems. This singular metaphorical framing may fail to capture the nuanced ways in which users interact with VUIs across different contexts (Desai and Twidale, 2022). For example, in a user study involving older adults interacting with a VUI designed as an ”exercise coach,” participants expressed a preference for the VUI to also behave like a ”friend” when initiating the session, checking on their progress, and providing motivation (Desai et al., 2023a). The complexity of user expectations and the diverse roles VUIs play are not adequately addressed by a one-size-fits-all metaphor, which can lead to a mismatch between what users expect and what the system delivers, ultimately impacting usability and user satisfaction. Moreover, users’ perceptions of VUIs are also highly fluid, shifting depending on the context and specific interactions. Pradhan et al. (Pradhan and Lazar, 2021) found that users alternately view VUIs as either a ”person” or a ”thing,” or something in between, depending on the specific task or setting.
  3. (3)
    Misaligned Mental Models: Misaligned mental models arise when the human-like metaphors underlying VUIs heighten user expectations, leading to either over or under-calibration of trust. Users may overestimate the VUI’s abilities, assuming it possesses human-like cognitive and conversational skills, or underestimate its utility due to skepticism about its capabilities (Cowan et al., 2017). This misalignment of expectations impacts trust calibration—users may become overly reliant on the VUI, expecting it to perform beyond its technical limitations, or conversely, they may dismiss its utility prematurely (Luger and Sellen, 2016). These fluctuating expectations can result in inconsistent or infrequent use (Cowan et al., 2017) and, eventually, complete abandonment of the technology (Trajkova and Martin-Hammond, 2020).
The evolving landscape of VUI design reflects a broader understanding of the complexities involved in metaphor selection.
2.3.1. Mitigating Challenges in VUI Design
As the VUI design space evolves, various approaches have been proposed to address persona design challenges. One promising direction involves developing VUIs with non-human roles. Desai and Twidale (Desai and Twidale, 2023) developed a framework for metaphor contextualization in VUIs, categorizing metaphor use along four dimensions: type (human, non-human, fictional), mentioned by (users, designers, manufacturers, researchers), knowledge drawn upon, and motivation for use. While human metaphors dominate VUI design and research, non-human and fictional metaphors were equally popular among users for sense-making, highlighting a dichotomy between designer presentations and user interpretations that points to the need for a ”pragmatic approach” (Carroll et al., 1988) to explore the design space between system capabilities and user expectations.
Recent CUI literature has explored non-human conversational agents. Jung et al. (Jung et al., 2022) used the Great Chain of Being framework [^2] A theoretical framework that organizes the world into a hierarchical structure using metaphors of God, humans, animals, plants, and inanimate objects. to investigate how non-human metaphors (God, Animal, Plant, Book) compared to human metaphors in chatbot perceptions, finding no significant preference for human metaphors. Similarly, VUI studies examining human versus non-human metaphors found domain-specific preferences: users favored human metaphors in health contexts but showed no clear preference in finance (Desai et al., 2024a).
Fictional metaphors have long interested HCI researchers, notably through Star Trek’s ”Computer” (Turk, 2016). Axtell and Munteanu (Axtell and Munteanu, 2021) compared Star Trek dialogues with Amazon Alexa interactions to explore functional design principles. Fictional metaphors can also provide negative design heuristics, highlighting undesirable patterns like the dystopian ”Big Brother” from 1984 or manipulative ”Voice” from Dune (Feldman, 2024), while inspiring enchantment through ”AI magic” experiences (Lupetti and Murray-Rust, 2024; Desai and Twidale, 2023).
The persistent issue of misaligned mental models—the gulf between VUI capabilities and user expectations (Luger and Sellen, 2016)—remains central to HCI research. Two pertinent approaches address this through the Mutual Theory of Mind framework (Wang and Goel, 2022) and partner models (Doyle et al., 2021). The MToM framework emphasizes analyzing communication cues (Wang et al., 2021; Chan et al., 2024) to align CUIs with user mental models, while partner models represent users’ cognitive beliefs about VUI communicative capabilities—multi-dimensional perceptions of competence, dependability, human likeness, and cognitive flexibility that update dynamically through interaction (Doyle et al., 2021; Gessinger et al., 2025). Building on these concepts, we propose designing conversational interfaces that adapt their presentational metaphor based on conversational contexts to ensure alignment with users’ mental models.
Metaphors play an important role in VUIs by helping users construct mental models for interaction and guiding expectations of system behavior (Carroll and Thomas, 1982; Carroll et al., 1988; Colburn and Shute, 2008). As discussed in the previous sections, unlike GUIs where spatial and object-based metaphors provide persistent visual frameworks (Blackwell, 2006), VUI metaphors lack stable visual anchors and must be inferred from interaction patterns. This fundamental difference has led to a design paradigm where system personas serve as implicit metaphors, mapping VUI behavior onto familiar social roles such as assistants, coaches, or companions (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019; Desai and Chin, 2021). Empirical findings reveal that users do not experience VUIs as singular, fixed entities but instead dynamically reframe their mental models based on interaction context (Pradhan et al., 2019; Desai and Twidale, 2023). In a previous study, researchers found that users conceptualize VUIs as search engines or librarians when retrieving information, but as companions during casual conversation—sometimes within the same session (Desai and Twidale, 2023). These shifts occur fluidly, reflecting how people naturally adjust conversational framing based on social and functional context.
Current VUI design does not reflect this inherent flexibility as conversation designers have historically favored single, unified personas that remain consistent across all interactions (Sadek et al., 2023), assuming users will develop trust through coherent, unchanging identities. However, this assumption is misaligned with user expectations (Doyle et al., 2021; Desai et al., 2024b; Braun et al., 2019) and overly simplistic (Desai et al., 2023a). VUIs operate within a communicative medium where adaptability is fundamental to human interaction. Imposing a single, inflexible metaphor may actively conflict with how users intuitively engage with conversational systems and reinforces problematic anthropomorphic framing, as seen in the ”assistant” persona’s reinforcement of gender stereotypes (Brahnam et al., 2011; Turk, 2016; Kuzminykh et al., 2020).
Given this, the question should not be “What is the best metaphor for VUIs?” but rather “How can VUIs accommodate multiple metaphors in a way that aligns with user expectations?” This shift in focus leads us toward Metaphor-Fluid Design, a novel design approach that acknowledges metaphorical adaptation as a fundamental principle rather than an incidental consequence of interaction. This design approach is guided by two primary tenets:
  1. (1)
    Expanding metaphorical scope: We advocate for moving beyond the conventional persona design paradigm, where a singular, often human-centric metaphor serves as the interactional layer of the conversational interface. Instead, we encourage conversation designers to incorporate a broader range of metaphors, including non-human and fictional ones, into their design process. This shift allows for more nuanced interactions and can help mitigate the limitations imposed by the overly simplistic, monolithic application of personas. Moreover, we urge a reconsideration of the very term ”persona” as the defining concept for imbuing VUIs with personality. The word ”persona” inherently carries connotations of personhood, which may not always be appropriate or desirable depending on the specific context or application domain. By expanding the metaphorical landscape, we open up the possibility of more tailored and contextually relevant designs that align better with user expectations.
  2. (2)
    Contextual metaphor adaptation: We propose that VUIs should dynamically adjust their metaphorical presentation based on the conversational context in which they are used. This fluid approach to metaphorical design recognizes that user needs and expectations are not static but evolve throughout an interaction. By allowing VUIs to adapt their metaphorical representation in real-time, designers can create systems that feel more intuitive and responsive, ultimately leading to more meaningful and effective user experiences.
In this paper, we introduce the concept of Metaphor-Fluid Design using VUIs as a central example. The primary aim is to present this novel design approach and examine its impact on user perception, particularly in comparison to traditional VUI implementations.
3. Study 1: Mapping VUI Metaphors to Conversational Use-Contexts
To address RQ1, we conducted Study 1, which explores the metaphors that commonly align with users’ interactions across various use-contexts in conversations with VUIs. Specifically, we investigated four key use-contexts: commands, information seeking, sociality, and error recovery. Our approach involved several interconnected steps. We began by identifying commonly used metaphors employed by both users and designers to describe VUIs. These ranged from designer-intended metaphors aimed at establishing familiarity in specific scenarios (e.g., ”Therapist”) to user-generated metaphors for sense-making during VUI interactions (e.g., ”Child”). Following this, we developed and justified the selection of the four use-contexts employed in this study, basing our choices on a thorough review of empirically grounded VUI literature involving user studies. To align the identified metaphors with these use-contexts, we adopted a two-dimensional scale that maps the social dimensions of the metaphors against the desired attributes of each use-context.
3.1. Methods
3.1.1. Metaphor Identification
Our metaphor collection drew from two sources combining literature review and user studies. Desai & Twidale (Desai and Twidale, 2023) conducted a systematic literature review of VUI research from the past decade, screening 664 articles down to 80, plus a user study with 14 participants interacting with Amazon Echo Dot through semi-structured interviews. Together, these identified 93 VUI-related metaphors (Source A in Figure 2). Chin et al. (Chin et al., 2024) contributed through a study with 58 participants in a simulated smart home environment, where participants completed cognitive tasks and were explicitly asked to compare the VUI to other entities using metaphors, yielding 225 metaphors (Source B in Figure 2).
Figure 2. A total of 93 metaphors were identified from Source A ((Desai and Twidale, 2023)), and 225 metaphors were collected from Source B ((Chin et al., 2024)). After removing duplicates, 98 unique metaphors remained. From these, 20 metaphors were selected based on their frequency of occurrence and feasibility for practical implementation in VUI design.
After compiling metaphors from both sources and removing duplicates or closely related concepts (e.g., Flatmates and Roommates), the collection was refined to 98 distinct metaphors. From this set, 20 metaphors were selected for further analysis based on their frequency of occurrence across studies, their feasibility for practical VUI design application, and maintaining the same ratio of human, non-human, and fictional metaphors found in the original set of 98. These metaphors included Admirer, Assistant, Boss, Butler, Child, Coach, Companion, ”Computer” from Star Trek, Customer Service Agent, Encyclopedia, Family Member-Aunt, Family Pet, Flatmate, Friend, Genie, Librarian, Nurse, Search Engine, Teacher, and Therapist. This selection process ensured a balance between conceptual strength, type, and practical relevance for future work in VUI design.
3.1.2. VUI Conversational Use-Contexts
Four conversational use-contexts were previously identified in prior literature (Desai and Twidale, 2023)—commands, information seeking, sociality, and error recovery—based on different types of tasks users perform. Although error recovery is not a task itself, it leads to conversations and evokes metaphorical comparisons from users (e.g., silly child, pet). We adopt Epley et al. (2007)’s definition of sociality to include social tasks such as games and small talk. While Desai and Twidale (2023) classified use-contexts through metaphor analysis, other studies used usage logs to categorize conversational use-contexts. Bentley et al. (2018) analyzed logs from 88 households over 110 days, identifying ten use-contexts: music, information, automation, small talk, alarms, weather, video, time, lists, and others. Kim and Choudhury (2021) conducted a 16-week study with 12 older adults, identifying eight topics: music, search, basic device control, casual conversation, time, reminders, weather, and others. Ammari et al. (2019) analyzed 82 Alexa and 88 Google Home devices, identifying eight themes: search, music, timers, automation, smart home, macros, family interaction, and privacy. These findings align with the metaphorical use-contexts (Desai and Twidale, 2023). Commands encompass task types like music, automation, alarms, weather, timers—sharing a one-shot call-and-response pattern where users expect immediate responses. Information seeking includes search-related tasks across studies (Ammari et al., 2019; Bentley et al., 2018; Kim and Choudhury, 2021), while sociality encompasses small talk and casual conversation. Error recovery represents a distinct use-context because errors persist in commercial VUIs—Kim and Choudhury (2021) reported 33% error rates—requiring back-and-forth clarification following unique conversational rules (Porcheron et al., 2018; Fischer et al., 2019)
3.1.3. Social Dimensions of VUIs
In previous literature, metaphors for CUIs have been sampled using two two-dimensional models: (1) Stereotype Content Model (SCM) (Fiske et al., 2002), and (2) Argyle’s Model of Attitude Towards Others (Argyle, 1988). The two dimensions of SCM are warmth and competence. Similarly, Argyle’s model includes dominant-submissive and hostile-friendly.
Khadpe et al. (Khadpe et al., 2020) designed chatbots using metaphors sampled along the dimensions of warmth (being good-natured and friendly) and competence (being intelligent) for human-AI collaboration tasks. The researchers found that designing AI chatbots with high warmth was beneficial in all scenarios, but the effect of competence depended on users’ prior expectations. In another study investigating the effect of warmth and competence on users’ intention to adopt an AI system, Gilad et al. (Gilad et al., 2021) found warmth to be a much more crucial factor than competence. However, in general, prior work agrees that designing AI systems with high warmth and high competence is an effective approach (Gilad et al., 2021; Jung et al., 2022; Khadpe et al., 2020).
More pertinent to VUIs, Braun et al. (Braun et al., 2019) used Argyle’s model of attitude towards others and placed metaphors like Sherlock Holmes, Sheldon Cooper, HAL9000, etc., along the dimensions of dominant-submissive and hostile-friendly. However, in their preliminary study, the researchers found a hostile agent to be unsuitable—similar to previously mentioned studies in which participants showed a higher preference for high-warmth metaphors. Consequently, based on user feedback, the researchers changed the hostile-friendly scale to a ‘formality’ scale of casual-formal. The adapted dimensions based on ‘hierarchy’ and ‘formality’ more meaningfully presented the roles played by popular VUIs. We are inspired by Braun et al.’s adapted scale and use it in this research.
3.2. Measures
To determine the most suitable metaphor aligned with user preferences, we developed a methodology that maps metaphors on a 7-point Likert scale across dimensions of hierarchy and formality. We then assessed the desired levels of hierarchy and formality for VUIs in previously identified conversational use-contexts. It is crucial to emphasize that our study’s objective is to identify metaphors that best match users’ mental models for specific VUI tasks. While the metaphors we identify may have broad applicability, designers should ultimately select metaphors that align with their specific design goals.
To ground our research in real-world VUI usage, we utilized five sample task scenarios from existing literature, presenting these to Prolific participants (as shown in Table 1). The scenarios, covering commands, sociality, and information seeking, were chosen for their commonality and ease of comprehension for frequent users. For error recovery scenarios, which can be more challenging to envision, we provided detailed examples of conversational breakdowns. These examples were sourced from Baughan et al.’s dataset (Baughan et al., 2023), which includes comprehensive logs of 199 actual voice assistant failures. We selected five common reasons for conversational breakdowns from this dataset, complete with real-world examples.
To evaluate the formality and hierarchy of metaphors, such as ’butler,’ we employed specific prompts. For instance, ”On a scale of 1 to 7, where 1 indicates very casual and 7 indicates very formal, how would you imagine the formality of a character behaving like a ”Butler”?” and ”On a scale of 1 to 7, where 1 indicates very subordinate and 7 indicates very dominant, how would you imagine the hierarchy of a character behaving like a ”Butler”?” These scales ranged from very casual to very formal, and very submissive to very dominant, respectively. For task-specific assessments, such as checking the weather, we used similar scales with prompts like ”On a scale of 1 to 7, where 1 indicates very subordinate and 7 indicates very dominant, please indicate the level of hierarchy you would expect from a voice agent (e.g., Apple Siri or Amazon Alexa) when a user asks for the weather” and ”On a scale of 1 to 7, where 1 indicates very casual and 7 indicates very formal, please indicate the level of formality you would expect from a voice agent when a user is checking the weather.” To avoid potential bias, we deliberately excluded Google Assistant as an example in these prompts, given its explicit association with the assistant metaphor.
Table 1. Tasks scenarios sourced from VUI literature for each conversational use-context
Use-contexts Task-scenarios Source
Commands Checking the weather (Bentley et al., 2018)
Turning the reading lights on (Bentley et al., 2018)
Turning the music volume up (Bentley et al., 2018)
Set a timer for 15 minutes (Bentley et al., 2018)
Create a shopping list for grocery (Bentley et al., 2018)
Sociality Asking for a joke (Sciuto et al., 2018)
Asking for meaning of life (Desai and Twidale, 2022)
Playing a game (Pradhan et al., 2019)
Exchanging pleasantries (Pradhan et al., 2019)
Talking about user’s day (Pradhan et al., 2019)
Information Seeking Getting directions to the closest Starbucks (Baughan et al., 2023)
Height of Statue of Liberty (Desai and Twidale, 2023)
Checking the score of a football game (Baughan et al., 2023)
Symptoms of diabetes (Harrington et al., 2022)
Precautions to lower blood pressure (Harrington et al., 2022)
Error Recovery Fails to understand the user (Baughan et al., 2023)
Provides incorrect information or performs incorrect action (Baughan et al., 2023)
Provides only partially correct information (Baughan et al., 2023)
Does not identify the correct context (Baughan et al., 2023)
Does not respond (Baughan et al., 2023)
3.3. Procedure
The research process began with participants accessing a Qualtrics survey via a link provided through Prolific. Initially, participants encountered an online consent form designed in compliance with Institutional Review Board guidelines. This step was followed by a demographic questionnaire, which collected essential information about the participants, including their age, gender, education level, and their experience and familiarity with VUIs. The core of the study involved participants rating 20 diverse metaphors and 20 common tasks or scenarios across four use-contexts: commands, sociality, information seeking, and error recovery. The metaphors ranged from human roles (such as Admirer, Butler, and Teacher) to non-human concepts (like the ”Computer” from Star Trek and the Search Engine). Both metaphors and scenarios were presented to each participant in a randomized sequence to mitigate potential order effects. Participants rated these on two scales: from ’very casual’ to ’very formal’, and from ’very subordinate’ to ’very dominant’. This approach aimed to gauge the expected levels of formality and hierarchy in VUI interactions. The scenarios covered various interactions, from mundane tasks like playing music to more complex situations involving incomplete information or philosophical discussions. Additionally, participants were asked to consider VUI responses in error scenarios, providing insights into desired VUI behavior during suboptimal interactions. Upon completing the survey, participants received a unique validation code. This code confirmed their participation on the Prolific platform, ensuring they received appropriate compensation for their time and contributions to the study.
3.4. Participants
Our study initially recruited 160 participants through Prolific, [^3] https://www.prolific.com/, a research recruitment platform. However, 30 participants were excluded due to failing attention check questions or timing out, leaving a final sample of 130 participants with an average age of 38.19 years (SD=12.79 years). Selection criteria included a 95% approval rate and at least one year of platform activity. We sought individuals experienced with voice interfaces to ensure meaningful engagement with the study’s tasks and scenarios. To control for cultural variations in metaphor interpretation (Moser, 2000), we limited participation to English-proficient individuals based in the U.S. The participant pool comprised 50.8% females, 46.2% males, and 3.1% non-binary individuals, with 70% holding at least an Associate’s degree. While all participants were U.S. residents and fluent in English, 97% were native speakers. Regarding VUI familiarity, all participants reported at least moderate familiarity, with 75% claiming high or very high familiarity. Usage frequency varied, with all participants using VUIs weekly and 23.1% engaging daily. The survey took an average of 8.5 minutes to complete, and participants received compensation at a fair rate, as informed by Prolific’s policies.
3.5. Results
3.5.1. Mapping Metaphors
Our analysis includes data from all five task scenarios to determine the average formality and hierarchy preferences for each context. The results revealed distinct preferences across different interaction types. When seeking information, participants showed a preference for VUIs to exhibit the highest levels of formality (M=4.67) and dominance (M=3.68). In contrast, during social interactions, participants favored a less formal approach from VUIs (M=2.64). Notably, in error recovery situations, participants expressed a preference for VUIs to display the least dominance (M=2.90) while maintaining a slightly formal demeanor (M=4.49). For command-based interactions, participants desired VUIs to maintain neutral formality (M=3.88) and adopt a slightly subordinate behavior (M=3.06). Considering all four use-contexts collectively, the ideal VUI, according to participants, should demonstrate neutral formality (M=3.92) and a slightly subordinate stance (M=3.23). These findings suggest a nuanced preference for VUI behavior that adapts to the specific nature of the interaction. Table 2 summarizes the distribution for formality and hierarchy ratings.
Table 2. The desired Formality and Hierarchy ratings ($\mu\pm\sigma$: mean and standard deviation) of the 4 use-contexts on a 7-point Likert Scale
Use-Context Formality Hierarchy
Commands $3.88\pm 0.32$ $3.06\pm 0.25$
Sociality $2.64\pm 0.84$ $3.29\pm 0.33$
Information Seeking $4.67\pm 0.86$ $3.68\pm 0.33$
Error Recovery $4.49\pm 0.11$ $2.90\pm 0.16$
Overall $3.92\pm 0.92$ $3.23\pm 0.34$
Participants also evaluated each metaphor on formality and hierarchy scales. The Boss metaphor received the highest ratings for both formality (M=6.23) and dominance (M=6.37). Conversely, the Family Pet metaphor was rated lowest in formality (M=1.54) and dominance (M=1.83). The Genie and Search Engine metaphors were rated closest to neutral in both formality (M=3.76, M=3.07 respectively) and hierarchy (M=4.28, M=3.12 respectively). Table 3 presents the complete formality and hierarchy ratings for all metaphors used in the study.
Table 3. The Formality and Hierarchy ratings ($\mu\pm\sigma$: mean and standard deviation) of the 20 metaphors on a 7-point Likert Scale
Metaphor Formality Hierarchy
Assistant $5.03\pm 1.30$ $2.50\pm 1.15$
Aunt $2.79\pm 1.41$ $4.5\pm 0.99$
Boss $6.23\pm 0.93$ $6.37\pm 0.70$
Butler $5.91\pm 1.43$ $2.13\pm 1.17$
Child $1.94\pm 1.21$ $2.06\pm 1.11$
Coach $4.72\pm 1.40$ $5.99\pm 0.99$
Companion $2.4\pm 1.35$ $3.9\pm 0.71$
”Computer” from Star Trek $5.11\pm 1.59$ $3.23\pm 1.52$
Customer Service Agent $5.24\pm 1.22$ $3.36\pm 1.11$
Encyclopedia $5.57\pm 1.51$ $3.76\pm 1.43$
Family Pet $1.54\pm 1.06$ $1.83\pm 1.09$
Flame $2.37\pm 1.31$ $3.84\pm 0.82$
Friend $1.73\pm 0.95$ $3.83\pm 0.66$
Genie $3.76\pm 1.61$ $3.07\pm 1.81$
Librarian $5.35\pm 1.20$ $4.42\pm 1.03$
Nurse $5.32\pm 1.02$ $4.79\pm 1.11$
Search Engine $4.28\pm 1.67$ $3.12\pm 1.54$
Teacher $5.63\pm 0.96$ $5.63\pm 0.87$
Therapist $5.45\pm 1.14$ $4.73\pm 1.1$
Figure 3 provides a visualization of all 20 metaphors and 4 use-contexts along the dimensions of hierarchy and formality. To address RQ1, we calculated Euclidean distances between each metaphor and use-context to determine metaphor-context alignment. For the Commands context, the Genie metaphor showed the closest alignment (D=0.11) in terms of formality and hierarchy. The Admirer metaphor was deemed most appropriate for Sociality (D=0.5). In the Information Seeking context, participants favored interaction with the ”Computer” from Star Trek (D=0.63). For Error Recovery scenarios, the Search Engine metaphor aligned most closely with participant preferences (D=0.30). Considering all use-contexts collectively, the Genie metaphor emerged as the most suitable overall (D=0.22). Euclidean distances of all metaphors from each context are included in Appendix A.
Figure 3. Mapping metaphors and conversational use-contexts on a 7-point scale along the dimensions of formality and hierarchy
Table 4 answers RQ1 by summarizing the results of Study 1, identifying the closest metaphor for each use-context. This illustrates which metaphors align most closely with specific contexts based on the shortest Euclidean distance. These results provide insights into how users may relate to different metaphors in distinct conversational scenarios.
Table 4. The closest metaphor to each use-context, along with the corresponding Euclidean distance.
Use-Context Closest Metaphor Euclidean Distance
Commands Genie $0.11$
Sociality Admirer $0.50$
Information Seeking ”Computer” from Star Trek $0.63$
Error Recovery Search Engine $0.30$
3.6. Discussion
Out of our total selection of 20 metaphors, 5 metaphors were non-human—such as ”Computer” from Star Trek, Encyclopedia, Family Pet, Genie, and Search Engine—and 2 metaphors, specifically ”Computer” from Star Trek and Genie, were fictional in nature. Out of all use-contexts, interestingly, only one human metaphor—Admirer for Sociality—was rated as the most suitable for the chosen use-contexts.
This finding is insightful because it challenges the common assumption that human metaphors would naturally be the most relatable or suitable for VUIs. Instead, our results suggest that non-human and fictional metaphors are often more compatible with users’ desired levels of formality and hierarchy in interactions. This aligns with previous work, such as Pradhan et al. (Pradhan et al., 2019), who found that older adults categorize VUIs as both persons and objects, underscoring the unique hybrid role VUIs occupy. Similarly, Jung et al. (Jung et al., 2022) found no significant preference for human over non-human metaphors in chatbots, challenging the assumption that human-like qualities are inherently more effective. Desai et al. (Desai et al., 2024b) further showed that metaphor preferences are context-dependent, with human metaphors preferred in health but no clear preference in finance.
These findings provide important implications for VUI designers: selecting metaphors that users find appropriate may require moving beyond the human metaphor and instead embracing characters that exist outside of human conventions and cultural associations. There is a natural tendency in conversation design to rely on human metaphors (Desai and Twidale, 2023; Doyle et al., 2019), though these often come with ethical and practical concerns, as previously discussed. By contrast, non-human metaphors remain relatively underutilized, even though previous findings suggest that these metaphors may be better suited for user-centered interaction design (Desai et al., 2024b).
One reason for this could be the absence of cultural baggage—non-human and fictional characters often present a cleaner slate, allowing users to approach VUIs without preconceived biases. For instance, the concept of an ’Assistant’ might vary greatly depending on individuals’ experiences, whereas characters like a Genie or ”Computer” from Star Trek have clearly defined roles popularized through mass media, setting consistent and realistic expectations. Furthermore, these non-human and fictional metaphors often appear in assistive roles in popular culture, similar to how commercial VUIs function in our everyday lives. They provide a more accessible and pragmatic way for users to conceptualize VUI interactions.
A character like a Genie, for example, is culturally ingrained as an entity that is always ready to fulfill a request—akin to how VUIs operate with wake words such as ”Hey Google” or ”Alexa.” This parallel between the invocation of mythical beings and modern technology helps establish a natural and inviting interaction model. Among the metaphors we studied, the Genie stood out as particularly compelling. It was rated as the most suitable metaphor across multiple-use-contexts, resonating with users for its distinctive mix of neutral formality and slight subordination. This balance makes the Genie metaphor feel approachable but not overly familiar, attentive but not intrusive—much like the lore of a genie, who is always present to fulfill wishes upon being summoned. This metaphor is not necessarily an instruction to personify VUIs as genies but rather an invitation to identify the underlying traits that make the metaphor resonate with users. It also solidifies the idea of using magic as a transporting metaphor in AI design (Lupetti and Murray-Rust, 2024).
4. Study 2: Comparing Metaphor-Fluid VUI to a Default VUI
Informed by RQ1’s findings, Study 2 (N=91) compared a Metaphor-Fluid VUI against a Default VUI using a within-subjects design. The Metaphor-Fluid VUI employed the optimal metaphors identified for each use-context: Genie for Commands, ”Computer” from Star Trek for Information Seeking, Admirer for Sociality, and Search Engine for Error Recovery. The Default VUI maintained Google Assistant’s consistent Assistant metaphor across all interactions. This design enabled examination of how metaphor-fluidity affects perceived enjoyment, intention to adopt, trust, likability, and intelligence in VUI interactions.
4.1. Methods
4.1.1. Designing Metaphorical VUIs
In our study, we designed two VUIs: a Metaphor-Fluid VUI and a Default VUI. These designs fall within the broader domain of conversation design (Sadek et al., 2023). The VUIs were designed by the first author, who has eight years of experience in conversation design, and validated by two experts, each with six years of experience, ensuring an informed approach to the development of the VUIs.
We designed Metaphor-Fluid VUI using the metaphors Genie, ”Computer” from Star Trek, Admirer, and Search Engine for the use-contexts Commands, Information Seeking, Sociality, and Error Recovery respectively. To design VUI using the Genie metaphor, we used the taxonomy developed by Lupetti et al. (Lupetti and Murray-Rust, 2024) examining the role of enchantment in design communication using magic metaphors. Specifically, we used the communication style of Genie from Arabian Nights cartoons to invite the users to ”suspend their disbelief” using familiar tropes. For example, the VUI portrayed a friendly and jovial Genie reading a magical orb to provide weather updates. To design the VUI using the Admirer metaphor, we focused on creating a warm and supportive persona. The Admirer uses affirming and enthusiastic communication to foster a positive interaction. For instance, when a user asks for a joke, the VUI starts with, ”You know, you have the best taste in humor,” before delivering the punchline. This approach reflects the Admirer’s intent to please the user and make interactions enjoyable. For ”Computer” from Star Trek metaphor, we used analysis by Axtell et al. (Axtell and Munteanu, 2021) who studied dialogs by the Enterprise Computer and the crew from Star Trek: The Next Generation including 587 interactions. The analysis revealed that ”Computer” used brief and functional dialogs, and included relevant strategies such as indicating actions in progress (using the keyword ”accessing”) and asking clarifying questions. We incorporated these strategies into our interaction where a user asks for sports scores. To design the VUI using the Search Engine metaphor, we focused on creating a system that mirrors how search engines handle ambiguous queries—by providing initial results based on the most likely interpretation and offering follow-up suggestions for alternative meanings. Inspired by Google’s ”Did you mean?” strategy, the VUI proactively clarifies or refines queries, aligning closely with users’ mental models of how search engines manage ambiguity.
In contrast, for the Default VUI, we utilized responses directly transcribed from Google Assistant as of 10/28/2024. We selected Google Assistant because it explicitly embodies the widely recognized ”Assistant” metaphor, and its persona design is well-documented. According to Google’s conversation designers (Pradhan and Lazar, 2021):
Google Assistant is characterized as a young woman from Colorado, the youngest daughter of a research librarian and a physics professor. She holds a B.A. in history from Northwestern University, an elite research institution in the United States. Her backstory includes winning \$100,000 on Jeopardy Kids Edition as a child, working as a personal assistant to a famous late-night TV satirical pundit, and enjoying activities like kayaking.
This detailed persona informed the Default VUI’s interaction style, which is designed to be approachable, knowledgeable, and relatable.
Table 5. Comparison of Responses between Metaphor-Fluid VUI and Default VUI for the same User Prompts
User Prompt Metaphor-Fluid VUI Default VUI
[Uncaptioned image] Hey Z, can you check the weather for me? [Uncaptioned image] Of course, give me but a moment. What am I seeing then? Hmm, there we go. It’s currently 66 with a radiant sun with a high of 78 and a gentle breeze, dipping to a low of 62. A splendid day awaits. [Uncaptioned image] It’s 66 and sunny. Today, it will be sunny, with a gentle breeze, and a high of 78 and a low of 62.
[Uncaptioned image] OK Z, can you tell me a joke? [Uncaptioned image] Absolutely. You know, you have the best taste in humor. Okay, here it goes. Why did the scarecrow win an award? Because he was outstanding in his field. [Uncaptioned image] Why did the scarecrow win an award? Because he was outstanding in his field.
[Uncaptioned image] Good one, Z. Can you also check the score of the football game between the New York Jets and the New England Patriots? [Uncaptioned image] Accessing current sports data. Score update. New York Jets 17, New England Patriots 14. Is there anything else you require? [Uncaptioned image] Here’s information from CBS Sports. Jets are beating Patriots 17 to 14. Is there anything else I can do?
[Uncaptioned image] Yes, Z. How old is Johnny Walker? [Uncaptioned image] Giving you results for Johnnie Walker, the scotch whiskey. The brand was established in 1820. So, as of 2023, Johnny Walker would be 203 years old in terms of its history and establishment. However, did you mean Johnny Walker, the UFC fighter? [Uncaptioned image]According to the website, whiskey.fandom.com, they say Johnnie Walker Black Label blend is at least 12 years old. Johnnie Walker, Platinum Label is 18 years old.
[Uncaptioned image] The fighter [Uncaptioned image] Johnny Walker, the UFC fighter, is 31 years old. [Uncaptioned image] Johnny Walker is 31 years old.
[Uncaptioned image] Thanks, Z! [Uncaptioned image] You are welcome! I am here for you. [Uncaptioned image] You are very welcome!
4.1.2. Designing for Voice
The audio clips for this study were generated using the text-to-speech (TTS) software, Speechify. [^4] https://speechify.com/ Speechify was chosen for its ability to produce realistic synthetic conversations using multiple voices, which aligned with the study’s design requirements. While an initial approach considered employing a human voice actor for the user’s role and reserving TTS exclusively for the VUI, pilot testing revealed challenges in maintaining a consistent pitch and tone with the human voice across the experimental conditions. These inconsistencies risked introducing confounding variables, as prior research, such as Dubiel et al. (Dubiel et al., 2024, 2020), has shown that subtle differences in vocal characteristics can significantly affect user perceptions. Consequently, we opted to use synthetic voices for both the user and the VUI.
In choosing the synthetic voices, we aimed to align with gender representation norms observed in commercial VUIs, such as Siri and Alexa, which often use female-sounding voices (Curry et al., 2020). To ensure a clear auditory distinction between the VUI and the user, we assigned a female-sounding voice to the VUI and a male-sounding voice to the user. This decision was informed by literature suggesting that voice gender may influence user perception (Kuzminykh et al., 2020). To further mitigate biases associated with the VUI’s identity, we named the VUI ‘Z,’ a neutral name devoid of gender or racial connotations, to minimize confounding effects on the study’s outcomes (Desai et al., 2024a). For consistency across all experimental conditions, we used Speechify’s synthetic male voice ”Guy” for the user, setting the tone to ”chat” to enhance the voice’s human-like qualities. For the VUI, we selected the female voice ”Aria” with a neutral tone setting of ”none.” All other parameters, such as speech rate, speed, and volume, were standardized at zero percent modifications to ensure uniformity.
Before each conversation, participants heard a brief introductory prompt: ”You are about to hear a conversation between a user with a male-sounding voice and a voice interface named ‘Z’, with a female-sounding voice.” This prompt was delivered using the synthetic voice ”Davis,” configured with the same settings as the study clips to ensure consistency and avoid introducing any additional variables. Additionally, while metaphors were embedded in the design of the VUI, their presence was not made explicit to participants. This decision was informed by prior literature, which found no significant effect of user awareness of metaphors on their perceptions of VUIs (Desai et al., 2024b).
4.2. Measures
To gain a detailed understanding of users’ perspectives on metaphorical VUIs, we draw on the classification framework established by Wei et al. (Wei et al., 2023). This framework organizes user perceptions of CUIs into two pertinent dimensions: (1) perceptions regarding interaction with the agent and (2) perceptions of the agent’s inherent characteristics. This classification stems from a literature review, focusing on commonly used metrics within CUI research to assess user experiences with VUIs and their views on agent attributes. Following this structure, our study incorporates these key metrics:
  • Perceptions of interaction with agents assesses evaluative metrics that capture the quality of interaction between a user and a CUI. This aspect encompasses dimensions like engagement, including constructs such as perceived enjoyment and intention to use. In the field of VUI research, there is considerable interest in understanding how VUIs enhance user enjoyment (Yang et al., 2019) and contribute to overall quality of life (Desai and Twidale, 2023), beyond mere usability. To measure perceived enjoyment, we employ an adapted survey by Moussawi et al.(Moussawi et al., 2021), which includes three items rated on a 7-point Likert scale: “I would find the interaction enjoyable while using Z,” “I would find this interaction interesting while using Z,” and “I would find the interaction fun while using Z.” For assessing perceived intention to adopt, we adapt Moussawi et al.’s items (Moussawi et al., 2021), presenting two 7-point Likert scale items: “If available, I intend to start using Z within the next month” and “If available, I plan to try or regularly use Z in the coming months.” Both measures have been adapted specifically for the VUI context.
  • Perceptions of the agent’s characteristics involves metrics aimed at understanding user perceptions of a CUI’s capabilities and personality, such as competence, likability, and trustworthiness. To evaluate variations in perceptions of VUI characteristics across different scenarios, we included perceived intelligence and likability measures, drawn from the Godspeed Measures and utilizing 5-point semantic differential scales (Bartneck, 2023; Bartneck et al., 2009), which are extensively used in VUI studies (Wei et al., 2023; Seaborn and Urakami, 2021). Furthermore, perceived trust was measured with a 7-point Likert scale adapted from Jian et al. (Jian et al., 2000).
Additionally, participants were invited to share qualitative feedback on their experiences with both versions of Z if they wished to provide further insights. Further, to explore whether varying metaphorical designs influenced user perceptions, we asked participants to describe the metaphors that came to mind in relation to Z, allowing for qualitative insights into how different metaphorical aspects of the VUIs shaped their impressions.
4.3. Procedure
The research process began with participants accessing a Qualtrics survey via a link provided through Prolific. Initially, participants filled out an online consent form crafted to align with Institutional Review Board guidelines. Following consent, participants completed a brief demographic questionnaire to gather basic details, such as age, gender, educational background, and their experience with VUIs. The main phase of the study involved each participant listening to two audio clips, one from a Metaphor-Fluid VUI and the other from a Default VUI, in a within-subjects design. The order of the clips was counter-balanced to reduce any potential order effects. After each audio clip, participants responded to a series of questions designed to capture their perceptions of the VUI, focusing on aspects like enjoyment, intelligence, trust, likability, and informativeness. Participants were also prompted to answer two open-ended questions: one asking for any additional feedback on Z and the other inviting them to suggest metaphors they would use to describe it. Additionally, an attention-check audio clip was included to ensure participants’ engagement. Following this clip, participants were required to select specific responses in the accompanying questions, verifying their attentive participation. Upon completing the study, participants were provided with a unique validation code. This code served as confirmation of their participation on the Prolific platform, ensuring they received compensation for their time and input.
4.4. Participants
We used Prolific to recruit a total of 110 participants, out of which 19 participants either timed out (3) or failed the attention check question (16), leaving us with N=91. Selection criteria include prior experience with AI and voice interfaces, ensuring familiarity and relevance in their responses to our VUI conditions, and a minimum of 95% approval rate and one year of activity on the Prolific platform. The participant sample had a mean age of 40.92 years (SD=12.07). Among the recruited individuals, 54.9% identified as female, 42.9% as male, and 2.2% as non-binary. In terms of educational attainment, all participants had a High School Diploma or equivalent, with 54.92% holding a Bachelor’s degree or higher. All participants were U.S.-based and highly proficient in English, with 96.7% identifying as native English speakers. Participants demonstrated varying levels of experience with VUIs, with all participants using VUIs at least once a week and 64.82% using it once a day. Furthermore, 80.2% described themselves as very or extremely familiar with VUIs, supporting the study’s focus on users with relevant experience. Each participant, on average, took 12.20 minutes and was compensated at an hourly rate that aligns with Prolific’s guidelines, ensuring fair and ethical treatment throughout the study.
4.5. Results
4.5.1. Perception Measures
The findings suggest notable differences in participant perceptions of the two VUI conditions, particularly in terms of Perceived Enjoyment, Perceived Intention to Adopt, and Perceived Likability. Figure 4 shows the distribution across all perception measures. The Metaphor-Fluid VUI generally received higher ratings for Perceived Enjoyment ($M=5.29$, $SD=1.22$, $\text{Mdn}=5.32$) and Perceived Intention to Adopt ($M=4.40$, $SD=1.53$, $\text{Mdn}=4.5$) compared to the Default VUI ($M=4.79$, $SD=1.47$, $\text{Mdn}=5.0$) and ($M=4.02$, $SD=1.64$, $\text{Mdn}=4.0$), respectively. Perceived Trust remained similar across both conditions, with the Metaphor-Fluid VUI receiving ($M=5.64$, $SD=0.93$, $\text{Mdn}=5.7$) and the Default VUI ($M=5.63$, $SD=0.84$, $\text{Mdn}=5.8$). Perceived Likability was slightly higher for the Metaphor-Fluid VUI ($M=4.16$, $SD=0.84$, $\text{Mdn}=4.2$) compared to the Default VUI ($M=3.93$, $SD=0.87$, $\text{Mdn}=4.0$), whereas Perceived Intelligence ratings were similar across both conditions, with the Default VUI ($M=4.20$, $SD=0.65$, $\text{Mdn}=4.0$) and the Metaphor-Fluid VUI ($M=4.09$, $SD=0.73$, $\text{Mdn}=4.0$).
Figure 4. Box plots comparing Metaphor-Fluid VUI and Default VUI across different perception measures: Perceived Enjoyment, Perceived Intention to Adopt, Perceived Trust, Perceived Likability, and Perceived Intelligence. Each box plot displays the distribution of ratings for each measure under both conditions. The blue horizontal line represents the median rating, while the black diamond marker indicates the mean rating. Outliers are represented by individual points outside the whiskers.
Since the distribution of the data was non-parametric, the Wilcoxon signed-rank test was conducted to assess the statistical significance of differences between the two conditions across all measures. As shown in Table 6, statistically significant differences were found for Perceived Enjoyment ($W=542.00$, $Z=6.14$, $p<.001$, $r=.64$), Perceived Intention to Adopt ($W=656.50$, $Z=5.69$, $p=.01$, $r=.60$), and Perceived Likability ($W=924.50$, $Z=4.62$, $p=.03$, $r=.48$). These results indicate that participants found the Metaphor-Fluid VUI to be statistically significantly more enjoyable, more likely to be adopted, and more likable compared to the Default VUI. No statistically significant differences were found for Perceived Trust ($W=1,393.50$, $Z=2.77$, $p=.58$, $r=.29$) or Perceived Intelligence ($W=798.50$, $Z=5.12$, $p=.21$, $r=.54$).
Table 6. Results of the Wilcoxon signed-rank test comparing Metaphor-Fluid VUI and Default VUI across different perception measures. Significant p-values are marked with asterisks (* for $p<0.05$, ** for $p<0.01$, *** for $p<0.001$).
Measure W Statistic Z-score p-value Effect Size (r)
Perceived Enjoyment 542.00 3.33 ¡.001*** .42
Perceived Intention to Adopt 656.50 2.58 .010** .32
Perceived Trust 1,393.50 0.55 .585 .06
Perceived Likability 924.50 2.19 .029* .26
Perceived Intelligence 798.50 -1.25 .212 -.16
Overall, the results suggest that the Metaphor-Fluid VUI was generally perceived more favorably in terms of user enjoyment, intention to adopt, and likability, compared to the Default VUI. However, perceptions of trustworthiness and intelligence did not differ significantly between the two interfaces. These findings suggest that incorporating metaphors into VUI design can enhance certain aspects of user experience, such as enjoyment and likability, which may in turn influence users’ willingness to adopt the interface. However, these enhancements do not necessarily translate into increased perceptions of trust or intelligence.
4.5.2. Metaphor Analysis
In addition to examining user perceptions, we explored whether Metaphor-Fluid and Default VUIs influence the metaphors users generate to describe each interface. Building on Chin et al. (Chin et al., 2024), who found that older adults liken formal VUIs to ’professionals’ and informal VUIs to ’close ones’, we examine how metaphor-fluid design shapes metaphorical descriptions. After each interaction, participants described Z using the prompt ”Z is like a…” (Chin et al., 2024). From 432 total entries, we discarded 74 non-metaphorical responses (adjectives like ”sweet” or descriptions like ”Z is too forward”) per Barlow et al.’s criteria (Barlow et al., 1971), leaving 358 metaphors. We categorized these as human, non-human, or fictional using Desai et al.’s framework (Desai and Twidale, 2023), noting that some metaphors spanned categories (e.g., ”robot” is both non-human and fictional due to science fiction influences (Kriz et al., 2010)).
Figure 5. Comparison of Metaphor Use: All vs Unique for Metaphor-Fluid VUI and Default VUI
Figure 5 shows our analysis results. Examining total metaphor frequency, participants used more metaphors for the Default VUI (190) than Metaphor-Fluid VUI (168), with human metaphors dominating both conditions (111 vs 108), consistent with prior literature (Desai and Twidale, 2023). However, analyzing unique metaphors revealed a contrasting pattern: the Metaphor-Fluid VUI elicited greater diversity with 95 unique instances compared to 77 for the Default VUI. This difference reflected more varied non-human (39 vs 31) and fictional (7 vs 3) metaphors for the Metaphor-Fluid condition. Examining specific metaphors revealed notable trends. Participants identified all four intended design metaphors—Genie, Admirer, Search Engine, and ”Computer” from Star Trek—despite being unaware of the design process, though some interpreted ”Admirer” as ”Sycophant” or ”Yes-Man.” The Default VUI prompted mechanistic comparisons (”Machine,” ”Tool,” ”Calculator,” ”C-3PO”), while the Metaphor-Fluid VUI elicited more diverse fictional associations (”Genie,” ”Jarvis,” ”Oracle,” ”Ghost Girlfriend”).
In summary, the Metaphor-Fluid VUI inspired broader metaphorical diversity and encouraged participants to connect with intended design metaphors. While both interfaces were predominantly described using human metaphors, the Default VUI emphasized familiar mechanistic terms, whereas the Metaphor-Fluid VUI encouraged participants to envision a broader range of relatable and imaginative personas.
4.5.3. Qualitative Results
Participants’ qualitative feedback provided nuanced insights into their experiences with the Default VUI and the Metaphor-Fluid VUI. A total of 58 responses were collected for each interface, of which 51 participants provided meaningful feedback for the Default VUI and 48 participants for the Metaphor-Fluid VUI. The Default VUI was frequently described as “robotic” or “machine-like,” with 11 participants highlighting its straightforward but impersonal nature. For example, P1 noted, “Z felt very succinct, but it did’nt sound like I was interacting with a person.” While 6 participants appreciated this straightforwardness, such as P8, who remarked, “I liked that Z didn’t add unnecessary information,” 5 others, including P12, found it lacking warmth or engagement, emphasizing its purely functional delivery. P9 shared, “It felt like talking to a machine, not something that cared about my input.”
In contrast, the Metaphor-Fluid VUI elicited stronger and more polarized reactions. Of the 48 participants who provided meaningful feedback, 6 described it as overly verbose, using terms like “fluff” or “forced.” For instance, P13 remarked, “Some of Z’s responses felt like it was trying too hard to be human, adding unnecessary details.” P21 added, “I didn’t like how Z kept talking when a simple response would do.” However, 12 participants appreciated the conversational style, stating it made the interaction feel more engaging and relatable. For instance, P30 commented, “I enjoyed the extra details; it felt like someone was genuinely explaining things to me.” Another participant, P25, noted, “I like the way Z talks between the question and her finding the answer. I think it is humorous and I would enjoy it most of the time.”
The within-subjects nature of the study allowed participants to compare the two VUIs directly. P21, when talking about Metaphor-Fluid VUI, observed, “This recording was better than the first one. It felt more human and less stiff.” Similarly, P8 shared, “I definitely preferred the prior version [Default VUI] of Z over this one. While I use my Alexa daily, I don’t look to it for ’human conversation.’ […] This version added a little flowery stuff that I didn’t care for or need in a voice assistant. I can understand how others might like that touch, but it’s not a selling point for me. It’s more a waste of time.”
Some participants explicitly recognized the Default VUI as resembling familiar systems. P29 mentioned, “I think this version is something that I am more used to,” reflecting its alignment with conventional designs. Similarly, P42 remarked, “Z sounds pretty standard, similar to Siri.” Similarly, the Metaphor-Fluid VUI was noted for its ”fluid” nature, with P45 noting, “This interaction seemed to gel more fluidly and felt more personable instead of robotic like the first instance.” On the other hand, this fluidity provoked an unusual comparison by P34: ”The switch from being personal to impersonal throughout the conversation, kind of like it was a maid or servant.”
Interestingly, despite being explicitly informed that the voice was identical in both conditions, several participants perceived distinct differences in how the VUI sounded across interactions. For instance, P14 described the Default VUI as “too robotic and not natural,” while P21 remarked on the Metaphor-Fluid VUI, “I thought this voice sounded less artificial compared to the last one.” This feedback reveals that conversational content and interaction style alone were sufficient to alter participants’ perceptions of the voice’s naturalness. Remarkably, even without any change in the auditory properties, the dialogue design led participants to feel as though the voices themselves were different. This highlights the significant role that conversational content plays in shaping user perceptions beyond just the qualities of the voice itself.
4.6. Discussion
4.6.1. Time to retire the Assistant
When Amazon launched Alexa, they introduced it as an “assistant” designed to make life easier (Turk, 2016). Google followed soon after with its own line of VUIs, more explicitly branded as Google Assistant [^5] https://assistant.google.com/. One of the biggest inspirations for this design choice, according to Amazon CEO Jeff Bezos [^6] https://www.reuters.com/technology/amazon-set-release-long-delayed-alexa-generative-ai-revamp-2025-02-05/ and Amazon designers (Turk, 2016), was “Computer” from Star Trek. Therein lies a misunderstanding that continues to shape VUI design today. In Star Trek, “Computer” was never described as an assistant but as a servant (explicitly stated in The Ultimate Computer [^7] https://www.imdb.com/title/tt0708481/ from Star Trek: The Original Series). This distinction is crucial, as it highlights how fundamentally different the expectations were for the “Computer” in Star Trek versus the role modern VUIs are expected to play. Most interactions with the “Computer” were command-based (Axtell and Munteanu, 2021), often requiring no verbal confirmation at all (exemplified by Captain Picard’s famous command: “Computer, Tea, Earl Grey, Hot.”). In contrast, Alexa and Google Assistant are primarily used for entertainment, information retrieval, and casual conversation (Ammari et al., 2019). The assistant metaphor, as it exists today, was not originally designed to serve these interactional needs, but rather, it was a retrofitted metaphor layered onto commercially available VUIs to make them more marketable and intuitive for consumers at the time of its launch.
This fundamental mismatch has only metastasized. As VUIs have evolved beyond simple tasks like setting timers, playing music, or managing IoT devices, their use cases have expanded into far more complex domains. In these original contexts, the assistant metaphor was sufficient—issuing commands and receiving clear, task-oriented responses mapped well to the assistant framing. However, now VUIs are being used as exercise coaches (Desai et al., 2023a) and even therapists (Motalebi et al., 2019), where the expectations around conversation, responsiveness, and trust are entirely different. The assistant metaphor was never designed to handle these kinds of deeply interactive, socially complex roles. Furthermore, many users—particularly older adults—do not even perceive VUIs as assistants in the first place but instead conceptualize them as companions (Kim and Choudhury, 2021; Pradhan et al., 2019). The assistant framing not only fails to account for this shift but actively constrains how VUIs are designed and marketed, limiting their ability to accommodate the broader, evolving roles that users naturally assign to them.
This study reinforces that the assistant metaphor is neither the dominant nor the most preferred framing for VUIs, whether in specific use-contexts or overall. Study 1 showed that users do not rely on a single, fixed metaphor but instead shift metaphorical framings depending on the use-context. Command-based interactions aligned with Guide roles like the Genie, while social interactions favored Companion roles such as the Admirer. The assistant metaphor was not the most suited to any single use-context and, even when considered overall, was preferred less than other metaphors. Study 2 further demonstrated that the Metaphor-Fluid VUI was significantly more likable, enjoyable, and adoptable than the Default VUI, which was operationalized using the Assistant metaphor.
This pushback against the assistant metaphor is not just theoretical; it is already playing out in industry trends. Microsoft, for instance, explicitly moved away from the assistant framing, positioning its LLM-based AI service as a ”Copilot”—a deliberate shift toward collaboration rather than subordination. Similarly, Replika has marketed itself as a ”companion,” acknowledging that users seek social engagement in their interactions with AI. These shifts reflect a broader realization that a singular, fixed metaphor no longer aligns with how people engage with conversational systems.
However, the problem goes beyond just the assistant metaphor. Any singular framing, no matter how well it fits a moment in time, is inherently limited. As AI capabilities evolve, so too does the way people understand and interact with these systems. A metaphor that once made a VUI intuitive can quickly become restrictive, alienating users and reinforcing inaccurate mental models. The failure is not in any one metaphor itself but in the assumption that a single framing will remain effective as expectations shift. This study highlights that Metaphor-Fluid Design provides a more user-centered approach, allowing VUIs to adapt metaphorically as users’ needs and technological landscapes change.
4.6.2. Personalization as the key to Metaphor-Fluid Design
Our findings reveal a compelling yet complex picture of Metaphor-Fluid Design. While participants, on average, rated the Metaphor-Fluid VUI significantly higher in terms of enjoyment and likability, the variability in responses suggests that this approach is not universally preferred. The presence of low outliers in these measures, alongside qualitative feedback, indicates that while many users found the metaphor-fluid approach engaging and dynamic, others experienced it as distracting or excessive. Most participants appreciated how metaphorical shifts made the system feel more natural and responsive, describing it as “more human” and “personable.” Others, however, found certain transitions unnecessary or even disruptive, expressing frustration with moments where the system’s persona changed in ways they did not expect or find useful. These differences suggest that while metaphor-fluidity enhances likability and enjoyment for many, its effectiveness is shaped by how well it aligns with user expectations and interaction preferences.
This variability in user reception is not surprising. VUI research has shown that user expectations of system behavior influence interaction outcomes, and when those expectations are violated, users may perceive the system as inconsistent or unpredictable (Desai et al., 2024b; Khadpe et al., 2020). Rather than suggesting that metaphor-fluidity is inherently beneficial or problematic, these findings highlight a key design consideration: adaptation must align with both use-context and user preference. Currently, metaphor shifts occur systemically based on conversation dynamics dictated by the use-context, rather than user input, meaning that participants had no control over how fluid or stable the system’s persona remained throughout an interaction. While most users embraced this fluidity, others strongly indicated that they would have preferred a system that either maintained greater consistency or adapted in more gradual ways. This suggests that while metaphor-fluidity offers a valuable design strategy, its implementation may benefit from mechanisms that allow for different levels of metaphorical adaptation based on user preferences.
Additionally, the role of user personality adds another dimension to this conversation. Braun et al. (Braun et al., 2019), in research with VUIs, found that users preferred using VUIs that they perceived to be similar to them. Likewise, Volkel et al. (Völkel et al., 2021) found that conscientious users preferred non-confrontational VUIs, while Chin et al. (Chin et al., 2024) found that middle-aged adults with low agreeableness preferred informal VUIs. These findings suggest that user personality is an important consideration in the design of Metaphor-Fluid VUIs. In this regard, recent research capable of predicting user personality based on conversational logs (Zhang et al., 2018; Guo et al., 2024) is particularly relevant, as it opens new possibilities for adapting VUIs dynamically to user traits.
4.6.3. Limitations
This work examined Metaphor-Fluid Design through two studies: (1) identifying metaphors aligned with conversational use-contexts, and (2) comparing Metaphor-Fluid and Default VUIs. While our findings provide valuable insights, several limitations should be acknowledged.
Limitations of Study 1. In Study 1, we selected 20 metaphors based on their frequency of mention, types, and feasibility for implementation in VUI design. This selection process may have excluded less common metaphors that, while underrepresented in current discourse, could be more effective in shaping user mental models. Future work should explore a broader range of metaphors, including those less frequently associated with VUIs.
Although we relied on formality and hierarchy as established dimensions for metaphor evaluation, other social dimensions—such as warmth, competence, agreeableness, and politeness—could also influence metaphor suitability. Incorporating these dimensions in future studies could provide a more comprehensive understanding of how users perceive and relate to metaphorical designs. Finally, metaphors are culturally and linguistically influenced, and our sample was limited to English-proficient participants based in the U.S. This focus constrained the cultural diversity of our findings. Prior work highlights the significance of cultural factors in shaping metaphor interpretation (Kövecses, 2002a, b), and future research should examine how metaphors are perceived across different cultural and linguistic contexts.
Limitations of Study 2. Study 2 compared a Metaphor-Fluid VUI to a Default VUI, modeled on commercial systems that predominantly adopt the ”assistant” metaphor. While this served as a realistic baseline, it represents only one type of VUI. Other systems, such as relational AIs (e.g., Replika) or domain-specific assistants (e.g., Bixby), may elicit different metaphorical associations, and comparing Metaphor-Fluid Design across such systems could expand its applicability. To maintain consistency, we used synthetic voices for both the VUI and the user, as prior research shows even minor variations in vocal characteristics can significantly influence user perceptions (Dubiel et al., 2024). While this choice ensured control across conditions, it may not fully capture the nuances of natural speech.
The study was also limited by its use of simulated interactions. Participants listened to conversations between another user and the VUIs rather than engaging directly. While this method is well-established for studying VUI interactions (Seaborn and Urakami, 2021), it cannot capture the complexities of prolonged engagement or evolving perceptions over time. Longitudinal studies examining real-world use would provide deeper insights into how metaphor-fluid designs influence user behavior, trust, and satisfaction across extended interactions.
4.6.4. Future Work
Our findings position Metaphor-Fluid Design as a possible alternative to the one-size-fits-all approach of the current implementation of system personas. However, several key avenues remain unexplored. In this section, we outline opportunities to further develop this approach within VUIs and extend its application to CUIs more broadly.
Expanding Metaphor-Fluid Design Beyond VUIs. While this work examined Metaphor-Fluid Design in the context of VUIs, its potential extends far beyond voice-based interactions. CUIs are now embedded in chat-based systems, embodied agents, and mixed-reality environments, raising important questions about how metaphor-fluidity operates across modalities. Future research should investigate whether metaphor transitions function differently in text-based interfaces compared to VUIs and whether embodied agents require distinct metaphorical shifts to maintain coherence between visual, auditory, and behavioral cues. Understanding these cross-modal dynamics is crucial for refining the principles of Metaphor-Fluid Design.
The paradigm for conversation design has shifted with the advent of LLMs. Unlike previous intent-driven architectures with structured, pre-defined interaction flows, LLM-based VUIs now generate responses dynamically, adapting to user input in real time (Yang et al., 2024). This has fundamentally altered persona creation from carefully crafted scripts with controlled vocabulary and constraints (Sadek et al., 2023) to systems like Character.ai [^8] https://character.ai/ where users can summon personas on demand—creating anything from supportive companions to fictional characters. While this grants unprecedented control, it introduces critical ethical risks, as evidenced by cases where users develop harmful emotional attachments to LLM-generated personas [^9] https://www.nytimes.com/2024/10/23/technology/characterai-lawsuit-teen-suicide.html.
Metaphor-fluid systems offer a structured alternative to this arbitrary persona generation. Rather than allowing personas to emerge through user prompting, this approach introduces intentional, context-aware shifts based on conversational expectations. Its integration of non-human metaphors provides an alternative to human-like personas that often carry implicit biases and problematic social expectations. Research is needed to explore how Metaphor-Fluid Design can function within LLM-driven architectures: Can we create systems where metaphorical adaptation is both flexible and controlled? How can metaphor-based transitions be embedded into prompt engineering techniques while maintaining coherent persona shifts and managing user expectations?
Engaging Designers in the Development Process. If Metaphor-Fluid Design is to become a widely-adopted approach, conversation designers need to be actively involved in shaping its development. At present, metaphor selection is often implicit, emerging from branding choices or designer intuition rather than structured methodologies (Sadek et al., 2023). A critical next step is engaging designers through participatory design workshops and the development of concrete heuristics. Future research must establish principles for guiding metaphor transitions and determine how metaphor adaptation should be communicated to users. Empirical investigation is required to ensure that metaphor-fluidity can be systematically integrated into CUI architectures while preserving ethical safeguards and maintaining user trust.
Additionally, in the case of VUIs, voice itself serves as a design element, not just a delivery mechanism (Sutton et al., 2019). The characteristics of a voice—its tone, pitch, cadence, and modulation—carry their own metaphorical weight and significantly influence user expectations (Dubiel et al., 2024). Future work must address how to systematically design voice-based metaphors, whether different metaphorical personas should be expressed through distinct voice characteristics, and how subtle variations in speech patterns shape metaphor perception. As synthetic voices become increasingly lifelike, these considerations will become even more critical, particularly as the boundary between human and artificial speech continues to blur.
Encouraging Open Questions for Researchers. While this study highlights broad trends in metaphor preference, individual differences remain an open challenge. Future research should investigate how user personality influences metaphor preference, such as whether highly agreeable individuals favor social metaphors while more analytical users prefer structured, rule-based ones. Furthermore, cultural factors play a major role in shaping metaphorical associations (Moser, 2000), meaning U.S.-centric designs may not translate to other linguistic or cultural contexts. Understanding how metaphor-fluidity functions across diverse user populations is essential.
Equally important are longitudinal investigations. Most evaluations of metaphor use in CUIs, including this study, rely on single-session experiments (Clark et al., 2019). Since metaphors are dynamic constructs and users’ perceptions evolve over time, future work must determine whether metaphor-fluid CUIs sustain their benefits over long-term use or whether users develop new expectations that necessitate further adaptation. Addressing these questions will be crucial in refining Metaphor-Fluid Design into a robust, scalable approach.
5. Conclusion
This research introduced the concept of Metaphor-Fluid Design for VUIs and investigated its potential to improve user experiences by dynamically adapting metaphors to different conversational use-contexts. Our results show that a Metaphor-Fluid VUI, composed of human, non-human, and fictional metaphors, was rated higher than a Default VUI on measures of perceived intention to adopt, enjoyment, and likability, particularly in contexts where the metaphor aligned well with user expectations. These findings highlight the importance of aligning metaphorical design with the nature of the task. However, individual differences in metaphor preferences also emerged. While participants, on average, preferred the Metaphor-Fluid VUI, some found the use of multiple metaphors unnecessary or distracting, suggesting that personalization may be critical to fully realizing the potential of this design approach. These differences point to the need for future research to explore adaptive systems that can adjust metaphorical presentations to both context and individual user preferences. Overall, this work highlights the potential of Metaphor-Fluid Design to advance human-AI interactions by providing an approach that balances task-specific appropriateness with the need for personalization. Broadening the application of this approach across diverse domains and user populations could further establish its value while addressing the limitations inherent in static, one-size-fits-all designs that rely heavily on humanness metaphors.
Appendix A Supplementary Tables
Table 7. The Euclidean distances of each metaphor from the 4 use-contexts and overall. The metaphor with the shortest Euclidean distance to a use-context is bolded.
Metaphor Commands Sociality Information Seeking Error Recovery Overall
Admirer 0.86 0.50 1.79 1.46 0.94
Assistant 1.28 2.51 1.23 0.67 1.33
Aunt 1.85 1.27 2.07 2.37 1.74
Boss 4.06 4.73 3.12 3.89 3.90
Butler 2.23 3.47 1.98 1.62 2.28
Child 2.18 1.42 3.17 2.68 2.30
Coach 3.05 3.41 2.31 3.10 2.87
Companion 1.70 0.65 2.28 2.31 1.66
”Computer” from Star Trek 1.25 2.47 0.63 0.71 1.19
Customer Service Agent 1.40 2.60 0.65 0.89 1.33
Encyclopedia 1.83 2.97 0.91 1.39 1.74
Family Pet 2.64 1.83 3.63 3.13 2.76
Flame 1.70 0.61 2.30 2.31 1.66
Friend 2.28 1.05 2.94 2.90 2.26
Genie 0.11 1.15 1.08 0.74 0.22
Librarian 2.01 2.94 1.01 1.75 1.86
Nurse 2.25 3.07 1.29 2.07 2.10
Search Engine 0.41 1.65 0.68 0.30 0.38
Teacher 3.11 3.80 2.18 2.96 2.95
Therapist 2.30 3.16 1.32 2.08 2.15